# Automated Knowledge Component Generation and Knowledge Tracing for Coding Problems

### v2

@@ 2.2 Improving Knowledge Tracing via LLM-generated KCs
Predictions To use LLMs to predict the student response to the next problem, we need to connect student KC knowledge with the textual input space of LLMs. Therefore, following previous work Fernandez et al. (2024); Liu et al. (2023), we transform KC mastery levels into soft text tokens, i.e.,
@@
where $\text{emb}^{\text{true}}$ and $\text{emb}^{\text{{false}}}$ are the embeddings of the text tokens “true”, and “false”, respectively. In other words, we use student KC mastery levels $m_{t}^{j}$ to combine two hard, discrete text tokens (“true” and “false”) into a differentiable soft token $s_{t}^{j}$, to enable the flow of gradients during training. We pass this student knowledge information using the input format of KC $1$: <$w^{1}$>. The student’s mastery level on <$w^{1}$> is: $s_{t}^{1}$. $\ldots$ into the LLM for prediction tasks.
Knowledge-Guided Response Prediction We construct our LLM prompt for the next response prediction by including both 1) the textual statement of the next problem and 2) student mastery levels on the KCs associated with the problem, as question: $p_{t}$. <KCs with student mastery levels>.
To predict the binary-valued correctness of the next student response, we average the hidden states of the last layer of Llama 3 that correspond to only the input (knowledge-guided prompt) to obtain a representation $r$, transformed for correctness prediction using a linear transformation matrix $W_{p}$ and a sigmoid function, given by $\hat{a}_{t+1}=\sigma(W_{p}\cdot r)$. We minimize the binary cross entropy (BCE) loss (for one response):
@@
To predict student code, we feed the knowledge-guided prompt into Llama 3 to generate the predicted code $\hat{c}$ token-by-token. We minimize
@@
Promoting Interpretability To promote interpretability of the student KC knowledge parameters, we use a compensatory model Maier et al. (2021) and take the average of individual student KC mastery levels to obtain an overall mastery level $\hat{y}_{t+1}=\frac{1}{\sum_{k=1}^{K}\mathbb{I}(w_{k})}\sum_{k=1}^{K}m_{t}^{k}\cdot\mathbb{I}(w_{k})$, where the indicator function $\mathbb{I}(w_{k})$ is $1$ if the KC $w_{k}$ is associated with the problem, and 00 otherwise. Empirically, we found that averaging over KC masteries performed better than taking a product over them, consistent with findings in prior work (Maier et al., 2021). We then minimize the BCE loss between this overall KC mastery level for this problem and its binary-valued correctness,
@@
Multi-task Learning Objective Following previous work Duan et al. (2025) showing multiple objectives in KT are mutually beneficial to each other, our final multi-task training objective minimizes a combination of all three losses together, with a balancing parameter $\lambda\in[0,1]$ controlling the importance of the losses, as
@@
3 Experimental Evaluation
We now detail our quantitative experimental settings to evaluate KCGen-KT on future student code submission correctness prediction.
Dataset Details The CodeWorkout DataShop (2021) dataset was first used in the Second CSEDM Data Challenge Challenge Organizers (2021) and contains actual open-ended code submissions from real students, collected from an introductory Java programming course, together with problem textual statements and human-written KC tags (estimated programming concepts) on each problem. In total, there are $246$ students attempting $50$ problems covering various programming concepts including conditionals, and loops, among others. Following prior work Shi et al. (2022), we only analyze students’ first submissions to each problem, leading to a total of $10,834$ code submissions.
Metrics For the binary-valued correctness prediction task, following Shi et al. (2022), we use standard metrics such as AUC, accuracy, and F1 score. For the student code prediction task, following Liu et al. (2022), we measure the similarity between generated student code and ground-truth student code using CodeBLEU Ren et al. (2020), a variant of the classic text similarity metric BLEU Papineni et al. (2002). This metric is customized for code and measures both syntactic and semantic similarity between two pieces of code.
Baselines In terms of KCs, we compare our generated KCs against human-written KCs that are available in the CodeWorkout dataset. We test a version of KCGen-KT by replacing our LLM-generated KCs with human-written KCs and keeping the KT method unchanged, which we refer to as KCGen-KT (Human-written KCs). In terms of KT methods, we adapt Test case-Informed Knowledge Tracing for Open-ended Coding (TIKTOC) Duan et al. (2025), a recent, strong KT method for programming, as the main baseline. TIKTOC also uses Llama 3 as the backbone and a multi-task learning setup to jointly predict the exact code token-by-token and whether it passes each test case. We slightly modify it for our KT task, replacing test case prediction with overall code correctness prediction, by reducing the dimension of the prediction head from the number of test cases to one, for overall correctness prediction only. We refer to the resulting method as TIKTOC*. We also use Code-DKT, a popular KT method for programming that leverages the content of student code, to predict the overall correctness of student code submissions. As a sanity check, to estimate a lower bound of performance on our KT task thereby providing a sense of task difficulty, we include two simple baselines: Random, which simply predicts the overall binary-valued correctness of a student code randomly with equal probability, and Majority, which simply predicts the majority correctness label (incorrect) among students for each problem.
Model KT Correctness Pred. Code Pred.
@@ 3 Experimental Evaluation
Random 0.499 0.368 0.506 --
Majority 0.500 0.644 0.526 --
Code-DKT Shi et al. (2022) $0.766_{\pm 1.8\%}$ $0.672_{\pm 3.3\%}$ $0.724_{\pm 1.0\%}$-
TIKTOC* Duan et al. (2025) $0.788_{\pm 1.3\%}$ $0.666_{\pm 3.0\%}$ $0.726_{\pm 1.3\%}$ $0.507_{\pm 1.5\%}$
KCGen-KT(Human-written KCs) $\underline{0.797}_{\pm 1.6\%}$ $\underline{0.706}_{\pm 2.6\%}$ $\underline{0.727}_{\pm 2.0\%}$ $\underline{0.557}_{\pm 2.8\%}$
KCGen-KT(Generated KCs) $\textbf{0.816}_{\pm 1.2\%}$ $\textbf{0.727}_{\pm 2.7\%}$ $\textbf{0.746}_{\pm 1.2\%}$ $\textbf{0.580}_{\pm 1.8\%}$
Table 2: Comparing KCGen-KT against baselines on KT performance across all metrics. KCGen-KT, especially with LLM-generated KCs, outperforms other KT methods. Best performance is in bold and second best is underlined.
Experimental Setup For the KT method component of KCGen-KT as well as for all KT baselines, to ensure a fair comparison, we use the instruction-tuned version of Llama 3 Llama Team (2024) with $8$B parameters as the base LLM and a frozen ASTNN Zhang et al. (2019) as the code embedding model. See Appendix B for detailed parameter settings. We repeat our experiments across 5 random train-validation-test data splits.
4 Results, Analysis, and Discussion
@@
We now discuss our quantitative evaluation results and qualitatively analyze the estimated student KC mastery levels and predicted code. We also analyze the learning curves, conduct an ablation study, and investigate the characteristics of KCs across different levels of abstraction.
4.1 Quantitative Evaluation
KCGen-KT outperforms baselines Table 2 shows the average performance (and standard deviation) on our two KT tasks: binary correctness prediction and student code generation for all methods. For KCGen-KT, we report results using the best-performing configuration, which generates $60$ KCs after clustering, based on $5$ student submissions per problem. We see that the Random and Majority baselines perform poorly, which suggests that the correctness prediction KT task is inherently difficult. Our proposed framework, KCGen-KT with either human-written or generated KCs, outperforms other strong KT methods that do not use KCs, including TITKOC* and Code-DKT. This observation suggests that for KT methods that use LLMs as the backbone, leveraging the semantic information in KC descriptions improves KT performance. More importantly, KCGen-KT with our generated KCs outperforms human-written KCs, by a consistent margin on both tasks, with statistical significance ($p<0.05$). This observation shows that high-quality KC descriptions and accurate tagging are key to improving downstream KT performance. The performance gap is more evident in code prediction, which shows that semantically informative KCs, as evident from Table 1, are especially important to LLMs in generative tasks.
Less Fine-grained KCs hurt performance To investigate the impact of KC granularity on model performance, we experiment with three levels of abstraction. We consider the 103 unique KCs generated by the first step of our pipeline, before clustering, as the most fine-grained (low-level) representation. We then apply the clustering algorithm with the number of clusters equal to 60 and 10 to get two other KC sets with medium and high abstraction levels. As the number of clusters decreases, the resulting KC sets become increasingly abstract, forming a hierarchy of representations. We evaluate KCGen-KT’s performance using these three KC sets. Table 3 shows that the highest abstraction level yields the lowest performance across all metrics on both tasks. In contrast, medium and low abstraction levels achieve comparable performance, which justifies our choice of using 60 KCs with medium-level abstraction. These results also suggest that overly abstract KCs may not pinpoint the necessary skills in a problem, underscoring the importance of having sufficient granularity in KCs for downstream student modeling tasks.
KC Abstraction Level AUC F1 Acc CodeBLEU
Low 0.815 0.726 0.737 0.572
Medium 0.816 0.727 0.746 0.580
High 0.794 0.683 0.708 0.557
@@ 4.1 Quantitative Evaluation
Table 3: Comparing different KC abstraction levels. SD omitted due to spatial constraints. Best performance is in bold and second best is underlined.
@@
Table 4: Example showing low student knowledge on relevant KCs map to specific errors in predicted student code.
Ablation Study We conduct an ablation study among all components of KCGen-KT. We find that including correct student submissions is crucial; removing it results in a noticeable performance decrease. We also explore the impact of using LLM-generated solutions and switching submissions to AST representation. See Appendix C for details.
4.2 Qualitative Evaluation

### v3

@@ 3.2.4. Soft KC Mastery Token Conversion
To use LLMs to predict the student response to the next problem, we need to connect student KC knowledge with the textual input space of LLMs. Therefore, following previous work (Fernandez et al., 2024; Liu et al., 2023), we transform KC mastery levels into soft text tokens, i.e.,
@@
where $\text{emb}^{\text{true}}$ and $\text{emb}^{\text{{false}}}$ are the embeddings of the text tokens “true”, and “false”, respectively. In other words, we use student KC mastery levels $m_{t}^{j}$ to combine two hard, discrete text tokens (“true” and “false”) into a differentiable soft token $s_{t}^{j}$, to enable the flow of gradients during training. We pass this student knowledge information using the input format of KC $1$: $w^{1}$>. The student’s mastery level on $w^{1}$> is: $s_{t}^{1}$. $\ldots$ into the LLM for prediction.
3.2.5. Knowledge-Guided Response Prediction
We construct our LLM prompt for the next response prediction by including both 1) the textual statement of the next problem and 2) student mastery levels on the KCs associated with the problem, as question: $p_{t}$. <KCs with student mastery levels>, as shown in Figure 2.
To predict the binary-valued correctness of the next student response, we average the hidden states of the last layer of Llama 3 that correspond to only the input (knowledge-guided prompt) to obtain a representation $r$, transformed for correctness prediction using a linear transformation matrix $W_{p}$ and a sigmoid function, given by $\hat{a}_{t+1}=\sigma(W_{p}\cdot r)$. We minimize the binary cross-entropy (BCE) loss, which for one response is given by:
@@ 3.2.5. Knowledge-Guided Response Prediction
To predict student code, we feed the knowledge-guided prompt into Llama 3 to generate the predicted code $\hat{c}$ token by token. We minimize
@@
3.2.6. Promoting Interpretability
To promote interpretability of the student KC knowledge parameters, we use a compensatory model (Maier et al., 2021) and take the average of individual student KC mastery levels to obtain an overall mastery level $\hat{y}_{t+1}=\frac{1}{\sum_{k=1}^{K}\mathbb{I}(w_{k})}\sum_{k=1}^{K}m_{t}^{k}\cdot\mathbb{I}(w_{k})$, where the indicator function $\mathbb{I}(w_{k})$ is $1$ if the KC $w_{k}$ is associated with the problem, and 0 otherwise. Empirically, we found that averaging over KC masteries performed better than taking the product over them, consistent with findings in prior work (Maier et al., 2021). We then minimize the BCE loss between the problem’s overall KC mastery and its binary correctness,
@@ 3.2.6. Promoting Interpretability
3.2.7. Multi-task Learning Objective
Following previous work (Duan et al., 2025) showing multiple objectives in KT are mutually beneficial to each other, our final multi-task training objective minimizes a combination of all three losses together, with a balancing parameter $\lambda\in[0,1]$ controlling the importance of the losses, as
@@ 3.2.7. Multi-task Learning Objective
Table 2. Comparing KCGen-KT against baselines on KT performance across all metrics. KCGen-KT, especially with LLM-generated KCs, outperforms other KT methods. Best performance is in bold and second best is underlined. $\bullet$ denotes statistically significant improvement over baselines. ($p<0.05$)
Dataset Model KT Correctness Pred. Code Pred.
@@
CodeWorkout (Java) Random 0.499 0.368 0.506 -
Majority 0.500 0.644 0.526 -
Code-DKT (Shi et al., 2022) $0.766_{\pm 0.018}$ $0.672_{\pm 0.033}$ $0.724_{\pm 0.010}$ -
TIKTOC* (Duan et al., 2025) $0.788_{\pm 0.013}$ $0.666_{\pm 0.030}$ $0.726_{\pm 0.013}$ $0.507_{\pm 0.015}$
KCGen-KT (Human-written KCs) $\underline{0.797}_{\pm 0.016}$ $\underline{0.706}_{\pm 0.026}$ $\underline{0.727}_{\pm 0.020}$ $\underline{0.557}_{\pm 0.028}$
KCGen-KT (Generated KCs) $\textbf{0.816}\textsuperscript{$\bullet$}_{\pm 0.012}$ $\textbf{0.727}\textsuperscript{$\bullet$}_{\pm 0.027}$ $\textbf{0.746}\textsuperscript{$\bullet$}_{\pm 0.012}$ $\textbf{0.580}\textsuperscript{$\bullet$}_{\pm 0.018}$
FalconCode (Python) Random 0.502 0.435 0.498 -
Majority 0.500 0.547 0.603 -
Code-DKT (Shi et al., 2022) $0.709_{\pm 0.013}$ $0.552_{\pm 0.021}$ $0.617_{\pm 0.019}$ -
TIKTOC* (Duan et al., 2025) $0.728_{\pm 0.011}$ $0.585_{\pm 0.019}$ $0.633_{\pm 0.006}$ $0.427_{\pm 0.010}$
KCGen-KT (Human-written KCs) $\underline{0.752}_{\pm 0.005}$ $\underline{0.599}_{\pm 0.013}$ $\underline{0.700}_{\pm 0.008}$ $\underline{0.473}_{\pm 0.007}$
KCGen-KT (Generated KCs) $\textbf{0.771}\textsuperscript{$\bullet$}_{\pm 0.011}$ $\textbf{0.645}\textsuperscript{$\bullet$}_{\pm 0.010}$ $\textbf{0.712}\textsuperscript{$\bullet$}_{\pm 0.006}$ $\textbf{0.498}\textsuperscript{$\bullet$}_{\pm 0.009}$
4. Experimental Evaluation
We now detail our quantitative experimental settings for evaluating KCGen-KT in predicting future student code and correctness.
4.1. Dataset Details
The CodeWorkout (DataShop, 2021) dataset was first used in the Second CSEDM Data Challenge (Challenge Organizers, 2021) and contains actual open-ended code submissions from real students, collected from an introductory Java programming course, together with problem textual statements and human-written KC tags (estimated programming concepts) from a fixed set of $18$ KCs on each problem. In total, there are $246$ students attempting $50$ problems covering various programming concepts. Following prior work (Shi et al., 2022), we only analyze students’ first submissions to each problem, leading to a total of 10,83410{,}834 code submissions.
The FalconCode  (de Freitas et al., 2023) dataset similarly consists of open-ended student submissions in an introductory Python programming course. The dataset provides problem statements and instructor-labeled KC tags from a predefined set of $20$ KCs. There are 3,2673{,}267 students attempting $157$ problems in both skill-based and lab-style formats. Similar to CodeWorkout, we analyze students’ first submission to each problem, resulting in a total of 28,61728{,}617 code submissions.
4.2. Baselines
In terms of KCs, we compare our generated KCs against human-written KCs that are available in both datasets. We test a version of KCGen-KT by replacing our LLM-generated KCs with human-written KCs and keeping the KT method unchanged, which we refer to as KCGen-KT (Human-written KCs). In terms of KT methods, we adapt Test case-Informed Knowledge Tracing for Open-ended Coding (TIKTOC) (Duan et al., 2025), a recent, strong KT method for programming, as the main baseline. TIKTOC also uses Llama 3 as the backbone and a multi-task learning setup to jointly predict the exact code token-by-token and whether it passes each test case. We slightly modify it for our KT task, replacing test case prediction with overall code correctness prediction, by reducing the dimension of the prediction head from the number of test cases to one, for overall correctness prediction only. We refer to the resulting method as TIKTOC*. We also use Code-DKT, a popular KT method for programming that leverages the content of student code, to predict the overall correctness of student code submissions. As a sanity check, to estimate a lower bound of performance on our KT task thereby providing a sense of task difficulty, we include two simple baselines: Random, which simply predicts the overall binary-valued correctness of a student code randomly with equal probability, and Majority, which simply predicts the majority correctness label (incorrect) among students for each problem.
4.3. Experimental Setup
We use repeated random subsampling with five splits, partitioning the data into training, validation, and test sets each time and averaging the results across runs. For the KT method component of KCGen-KT as well as for all KT baselines, to ensure a fair comparison, we use the instruction-tuned version of Llama 3 (Llama Team, 2024) with 88B parameters as the base LLM and a frozen ASTNN (Zhang et al., 2019) as the code embedding model. We use the Parameter Efficient Fine-Tuning (PEFT) library from HuggingFace (Wolf et al., 2020) to load Llama 3 and fine-tune it via Low-Rank Adaptation (LoRA) (Hu et al., 2022) ($\alpha=256$, $\text{rank}=128$, $\text{dropout}=0.05$) using 8-bit quantization. We use the AdamW (Loshchilov and Hutter, 2019) optimizer for LLM and $W_{m}$ parameter with a batch size of $32$ and the RMSprop optimizer for the LSTM, and perform a grid search to determine the optimal learning rate. In KCGen-KT, we set different learning rates for different model components: 1e51e-5 for Llama 3, 5e45e-4 for the LSTM model, and 1e41e-4 for the $W_{m}$ and $b_{m}$ parameters. KCGen-KT converges within $12$ training epochs on both datasets, with each epoch taking $80$ minutes on the CodeWorkout dataset and $300$ minutes on the larger FalconCode dataset when trained on an NVIDIA L40S 48GB GPU.
For the baseline model, TIKTOC*, which also uses an LLM as the backbone of the model, we use the same setup for a fair comparison: we use the same Llama model with PEFT to load it and fine-tune it with LoRA ($\alpha=256$, $\text{rank}=128$, $\text{dropout}=0.05$) using 8-bit quantization. The optimizers and learning rate used are the same as KCGen-KT, except the learning rate for the LSTM model is 1e41e-4. For all models that perform the code prediction task, we use greedy decoding to generate student codes.
4.4. Metrics
Following (Shi et al., 2022), we use AUC, accuracy, and F1 score to evaluate the correctness prediction performance. For the student code prediction task, following (Liu et al., 2022), we measure the similarity between generated student code and ground-truth student code using CodeBLEU (Ren et al., 2020), a variant of the classic text similarity metric BLEU (Papineni et al., 2002). This metric is customized for code and measures both syntactic and semantic similarity between two pieces of code.
5. Results, Analysis, and Discussion
@@
We now discuss our quantitative evaluation results and qualitatively analyze the estimated student KC mastery and predicted code. We also analyze the learning curves, conduct an ablation study, and investigate the patterns of KCs across different levels of abstraction.
5.1. Quantitative Evaluation
@@
Table 3. Comparing different KC abstraction levels across datasets. SD omitted for space. Best performance is in bold and second best is underlined.
Dataset Clusters (Lvl) AUC F1 Acc CodeBLEU
100 (Low) 0.815 0.726 0.737 0.572
75 (Low) 0.810 0.719 0.735 0.575
CodeWorkout 50 (Medium) 0.816 0.727 0.746 0.580
20 (High) 0.804 0.698 0.710 0.561
10 (High) 0.794 0.683 0.708 0.557
150 (Low) 0.763 0.628 0.675 0.494
120 (Low) 0.763 0.651 0.706 0.483
FalconCode 90 (Medium) 0.762 0.648 0.674 0.497
60 (Medium) 0.767 0.653 0.701 0.503
20 (High) 0.750 0.623 0.663 0.473
5.1.1. KCGen-KT outperforms baselines
Table 2 shows the average performance (and standard deviation) on our two KT tasks: binary correctness prediction and student code generation for all methods on both datasets. For KCGen-KT, we report results using the best-performing configuration for each dataset, which generates $50$ KCs after clustering for CodeWorkout and $60$ KCs after clustering for FalconCode, based on $5$ student submissions per problem. Across both datasets, we see that the Random and Majority baselines perform poorly, which suggests that the correctness prediction KT task is inherently difficult. Our proposed framework, KCGen-KT with either human-written or generated KCs, outperforms other strong KT methods that do not use KCs, including TITKOC* and Code-DKT. This observation suggests that for KT methods that use LLMs as the backbone, leveraging the semantic information in KC descriptions improves KT performance. More importantly, KCGen-KT with our generated KCs outperforms human-written KCs, by a consistent margin on both datasets and across both tasks, with statistical significance ($p<0.05$). This observation shows that high-quality KC descriptions and accurate tagging are key to improving downstream KT performance. The performance gap is more evident in code prediction, which shows that semantically informative KCs, as evident from Table 1, are especially important to LLMs in generative tasks.
5.1.2. Less Fine-grained KCs hurt performance
To investigate the impact of KC granularity on model performance, we experiment with multiple levels of abstraction across two datasets. For the CodeWorkout dataset, our pipeline first generates $103$ unique KCs, then applies the clustering algorithm with the number of clusters set to $100$, $75$, $50$, 2020, and $10$. For the FalconCode dataset, we begin with $161$ unique KCs and then cluster them into $150$, $120$, $90$, $60$, and $20$ clusters. As the number of clusters decreases, the resulting KC sets become increasingly abstract, forming a hierarchy of representations. We evaluate KCGen-KT’s performance under all abstraction levels. Table 3 shows that the highest abstraction levels consistently lead to the lowest performance across all metrics and datasets. In contrast, medium and low abstractions (e.g., 50 or 100 clusters for CodeWorkout, 60 or 120 for FalconCode) achieve comparable performance, which justifies our choice of using medium-level KCs in the main experiments. These results suggest that overly abstract KCs may fail to pinpoint the necessary skills in a problem, underscoring the importance of having sufficient granularity in KCs for downstream student modeling tasks.
Table 4. Ablation study of KCGen-KT on CodeWorkout.
Model AUC F1 Acc CodeBLEU
$\blacksquare$ KCGen-KT 0.812 0.723 0.724 0.569
$\square$ w/o Correct Solutions 0.789 0.674 0.704 0.529
$\square$ w/ Incorrect Solutions 0.773 0.651 0.700 0.516
$\square$ w/o In-context Examples 0.782 0.677 0.705 0.539
$\square$ Partial In-context Examples 0.810 0.723 0.717 0.562
$\square$ Student Code $\rightarrow$ AST 0.784 0.691 0.715 0.546
$\square$ Student Code $\rightarrow$ Generated 0.807 0.706 0.721 0.557
5.1.3. Ablation Study
Table 4 shows the results of the ablation study among all components of KCGen-KT on CodeWorkout. We chose CodeWorkout as the representative dataset due to its more balanced score distribution and cleaner, method-only code format. In contrast, some FalconCode problems include starter codes that may introduce confounding effects. For ablation settings requiring code submissions for KC generation, we report results using two code solutions since we note that LLM-generated code tends to follow similar problem-solving strategies; we adopt the most fine-grained KC abstraction (100 clusters) and manually verify the correctness of all generated solutions. We see that including correct submissions is crucial for KC generation; removing them results in a performance decrease on both KT tasks, indicating that the problem statement alone is insufficient to capture all necessary KCs. Similarly, in-context examples also play a critical role. Removing them results in generating overly abstract KCs, which are less effective for the KT tasks. We also test replacing the complete problem–KC mapping in the in-context example with only two randomly selected KCs, with the performance obtained being comparable to KCGen-KT, suggesting that minimal human involvement is required.
Table 5. Ablation study of KCGen-KT on CodeWorkout on different number of solution codes provided during KC Generation. Best performance in bold and second best is underlined.
No. of Solutions AUC F1 Acc CodeBLEU
1 0.804 0.713 0.705 0.563
2 0.812 0.723 0.724 0.569
5 0.815 0.726 0.737 0.572
7 0.812 0.726 0.727 0.573
10 0.816 0.727 0.725 0.566
Table 6. Example showing low student knowledge on relevant KCs map to specific errors in predicted student code.
@@ 5.1.3. Ablation Study
We further investigate incorporating both correct and incorrect solutions during KC generation in order to capture mistake-related KCs. However, this ablation hurts performance, likely because KCs derived from incorrect solutions encode errors or misconceptions, thereby introducing noise into the soft-token-based model. Replacing actual student submissions with LLM-generated submissions also results in a minor performance decrease, which suggests that actual student code is highly diverse and captures a more complete set of skills required in each problem. We also see that switching submissions to the abstract syntax tree (AST) representation as input to the LLM decreases performance, where we see the generated KCs tend to be less detailed. This result can be explained by LLMs not being heavily pre-trained on AST representations of code.
Finally, we explore the impact of the number of student submissions provided to the LLM on the quality of generated KCs and KT performance on CodeWorkout. For this ablation, we use the most fine-grained abstraction level. Table 5 shows the results, where we see that using fewer submissions, such as one and two, results in worse performance, while the performance increases and saturates after more than five submissions. This result can be explained by a smaller number of student submissions failing to capture diverse solution strategies for some problems, thus resulting in an incomplete KC set. As the number of submissions increase, the set of initial KCs before clustering more or less stays the same, and the performance mostly depends on the abstraction level instead.
5.2. Qualitative Evaluation
5.2.1. Case Study